Find an equation of the line having the given slope and containing the given point: m=3; (3,5)
Answer by mananth(16946)

m= 3

Plug value of the slope and point(3,) in
y=mx+b
5.00 = 3 * 3 +b
b= 5 -9

b= -4
So the equation will be
Y = 3 x + -4
m.ananth@hotmail.ca